40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6056     Entry time: Thu Dec 1 01:24:56 2011
Author: Jenne 
Type: Update 
Category: RF System 
Subject: EOM temp stabilization at PSL lab 

[Frank, Jenne]

Activities in a far, far away land called PSL lab...

EOM_Temp_stab_30Nov2011.png

We are looking at the RFAM from the detuned ACAV refl pd in the red trace.  Red trace is the demodulated RFAM output.  RCAV was locked, so on a ~min time scale the freq drift follows, so we stay detuned. 

Heater and temp sensor are taped to EOM, no foam box.

Around when the red trace starts, we turn on the heater to stabilize the temp.  After a while we reach the set point (no longer railing the heater), and start seeing temp stabilization.  The RFAM fluctuations clearly go down.  Neato. 

No calibrations, no RIN, no nothing.  Get over it.

Green trace is the DC level of a different PD, which should also be sensitive to RFAM.

This is using a fancy-pants temp controller chip that Frank found.  It works, so that's handy.

ELOG V3.1.3-